OpenStudy (wasiqss):

Calculate equation of rectangular hyberbola with its focal length (128)^1/2

OpenStudy (jamesj):

If xy = c is the equation of a rectangular hyperbola, what's the equation for the focal length?

OpenStudy (jamesj):

Once you know it, set it equal to (128)^1/2 and solve.

OpenStudy (wasiqss):

c^2=a^2+b^2

OpenStudy (wasiqss):

how did u drive xy=c

OpenStudy (wasiqss):

ok

OpenStudy (2bornot2b):

You need to know the eccentricity I think

OpenStudy (wasiqss):

yeh

OpenStudy (2bornot2b):

Is it given in the problem?

OpenStudy (wasiqss):

noo

OpenStudy (2bornot2b):

Ok then I typing the solution, with the letter e

OpenStudy (2bornot2b):

Focal length of a hyperbola is ae, which according to you is equal to sqrt(128), then I think a is equal to {sqrt(128)/e}. Since it is a rectangular hyperbola, a = b So the equation to the hyperbola is \[\frac{x^2}{\sqrt{128}/e}+\frac{y^2}{\sqrt{128}/e}=1\]

OpenStudy (2bornot2b):

..............where e is the eccentricity. I need others' comment on this question, to be able to confirm if I am right in saying that the data of e should have been supplied by the question setter to make this question complete.

OpenStudy (wasiqss):

i guess e shud b eliminated sum ow

OpenStudy (jamesj):

Wait. You asked about a hyperbola, not an ellipse. That is why I wrote down the equation of a hyperbola.

OpenStudy (wasiqss):

yeh i asked about hyperbola

OpenStudy (jamesj):

the problem is, the equation above is the equation of an ellipse.

OpenStudy (2bornot2b):

Oops! that equation with a -ve sign in the middle

OpenStudy (wasiqss):

which equation

OpenStudy (jamesj):

n any case, there are two standard form for a hyperbola. right ... the one I wrote down and then the one above corrected with a minus sign.

OpenStudy (2bornot2b):

Focal length of a hyperbola is ae, which according to you is equal to sqrt(128), then I think a is equal to {sqrt(128)/e}. Since it is a rectangular hyperbola, a = b So the equation to the hyperbola is \[\frac{x^2}{\sqrt{128}/e}-\frac{y^2}{\sqrt{128}/e}=1\]

OpenStudy (wasiqss):

i want e eliminated sum how, susanne give it a shot, n also check ur fan box